# Quantum State Preparation with Optimal T-Count

### v1

@@ 3 Quantum state preparation with optimal T-count
Our overall state preparation approach is similar to the one by Rosenthal [40]. Movtivated by the Aaronson-Kuperberg problem [2, 1], [40] focuses on the number of Boolean oracle calls needed for efficient state preparation. While the context is different, the ${T}$-count of the algorithm presented in [40] is near optimal. To obtain our Theorem 1.1, we combine Rosenthal’s algorithm with the crucial missing pieces Theorem 1.2 and Theorem 1.3. Along the way, we also simplify and improve some of Rosenthal’s analysis: our Lemma 3.2 simplifies and improves [40, Lemma 3.2], which was implicit in [25]; our Lemma 3.5 improves the query complexity in [40, Theorem 4.2].
We already know from Theorem 1.2 how to implement an arbitary diagonal unitary using the claimed number of ${T}$ gates and ancillas. For the state preparation task, this lets us specialize without loss of generality to the case where the target state has real amplitudes: after preparing a target state with real entries, we can apply a diagonal unitary to apply any desired complex phases.
@@
Let $|\psi\rangle$ be an arbitrary $n$-qubit state with real amplitudes. There exists an $n$-qubit state $|\phi\rangle$ with real amplitudes such that $\left\langle\phi\middle|\psi\right\rangle\geq\frac{1}{\sqrt{2}}$ and $|\phi\rangle=B_{2}{H}^{\otimes n}B_{1}{H}^{\otimes n}|0^{n}\rangle$ for some $n$-qubit Boolean phase oracles $B_{1}$ and $B_{2}$ (i.e., diagonal unitaries in which all diagonal entries are ±$\pm 1$).
@@
Let $B$ be an $n$-qubit Boolean phase oracle. Then $B$ can be prepared exactly by a Clifford+${T}$ circuit using $O\left(\sqrt{2^{n}}\right)$ ${T}$ gates and ancillary qubits.
As a result of Fact 3.3, we can prepare the approximation $|\phi\rangle$ from Lemma 3.2 using only $O\left(\sqrt{2^{n}}\right)$ ${T}$ gates and ancilla qubits. We see that Lemma 3.2 already establishes the special case of Theorem 1.1 where the error $\varepsilon$ is a sufficiently large constant.
@@
Using Theorem 1.2 to correct phases, we assume without loss of generality that the target state $|\psi\rangle$ has real amplitudes. Then it suffices to analyze the ${T}$-count and ancilla count for the circuit in Lemma 3.5. To this end, we analyze the components separately, each of which appears constant number of times in the whole circuit in Lemma 3.5. Assume $\varepsilon\leq 1/2$.
  • The reflection $2\left|0^{n+t}\right\rangle\left\langle 0^{n+t}\right|-I$ used in $R_{1}$ is a Boolean phase oracle and can be constructed by Fact 3.3 with $O\left(\sqrt{2^{n+t}}\right)=O\left(\sqrt{2^{n}\log(1/\varepsilon)}\right)$ ${T}$ gates and ancillas. [^13] This reflection can be implemented with much lower cost. We use Lemma 2.1 as a loose upper bound since it does not affect the asymptotic bound in the end.
@@
  • The Hadamard layer used in $V$ is Clifford and requires no ${T}$ gates or ancilla.
  • The Boolean phase oracle on $\mathcal{A}\mathcal{B}$ used in $V$ can be again prepared using Fact 3.3 with $O\left(\sqrt{2^{n+t}}\right)=O\left(\sqrt{2^{n}\log(1/\varepsilon)}\right)$ ${T}$ gates and ancillas.
  • The layer of single-qubit gates on $\mathcal{A}$ used in $V$ can be synthesized by Theorem 1.3 since $\mathcal{A}$ only has $t=\log\log(1/\varepsilon)+O(1)$ qubits. This uses $O\left(\log(1/\varepsilon)\right)$ ${T}$ gates and ancillas by Theorem 1.3. The approximation error does not blow up since this is only used a constant number of times in total.
@@
To get the coarse approximation of Lemma 3.2, we take our target state and apply a random diagonal unitary with ±$\pm 1$ entries followed by a binary Fourier transform (i.e., the Hadamard matrix). This does a reasonable job of flattening the amplitudes of the state, by which we mean that most of the amplitudes are now of similar magnitude. We then apply a diagonal unitary to make all the phases $+1$, which gives a state that has constant overlap with the uniform superposition $|+^{n}\rangle$. To implement this strategy we use the standard Khintchine inequality.
@@
After flattening, the state becomes essentially a uniform superposition with different phases. Moreover, in the case of interest where $|\psi\rangle$ is assumed to have real amplitudes, these phases are simply ±$\pm 1$. Then we can correct the Boolean phases and reverse the flattening procedure to obtain an approximation of the original state $|\psi\rangle$.
@@
Since $|\psi\rangle$ has real amplitudes, so does $|\widetilde{\psi}\rangle$. Let $B_{1}$ be the Boolean phase oracle such that $\langle x|B_{1}|x\rangle=\mathrm{sign}(\langle x|\widetilde{\psi}\rangle)$ for all $x\in\{0,1\}^{n}$.
@@
Let $\alpha\in(0,1]$ be a real number. For any $n$-qubit state $|\phi\rangle$, let $U_{\phi}$ be an $n$-qubit unitary that satisfies $\Re\langle\phi|U_{\phi}|0^{n}\rangle\geq\alpha$ [^14] For a complex number $v$, $\Re v$ equals its real part. Define
@@
Since $\Re\langle\phi|U_{\phi}|0^{n}\rangle\geq\alpha\geq\gamma$, we have
$\left\||\phi\rangle-\gamma\cdot U_{\phi}|0^{n}\rangle\right\|=\sqrt{1+\gamma^{2}-2\gamma\cdot\Re\langle\phi|U_{\phi}|0^{n}\rangle}\leq\sqrt{1-\gamma^{2}}=\beta.$ (22)
@@
$=\left\||\widetilde{\psi}_{j-1}\rangle\right\|^{2}+\gamma^{2}\beta^{2j-2}-2\gamma\beta^{j-1}\left\||\widetilde{\psi}_{j-1}\rangle\right\|\cdot\Re\langle\psi_{j-1}|U_{\psi_{j-1}}|0^{n}\rangle$ (26)
$\leq\left\||\widetilde{\psi}_{j-1}\rangle\right\|^{2}+\gamma^{2}\beta^{2j-2}-2\gamma^{2}\beta^{j-1}\left\||\widetilde{\psi}_{j-1}\rangle\right\|.$ (since $\Re\langle\psi_{j-1}|U_{\psi_{j-1}}|0^{n}\rangle\geq\gamma$)
@@
In light of Lemma 3.8, we set $\alpha=\frac{1}{\sqrt{2}}$ and design $U_{\phi}$ as the $B{H}B{H}$ circuit from Lemma 3.2. Note that all states used in Lemma 3.8 have real amplitudes since we start with $|\psi\rangle$ that has real amplitudes. Therefore circuits $U_{\phi}$ are all well-defined. ∎
@@
In this Section we describe the circuit that prepares an $n$-qubit state with optimal ${T}$-count, establishing Lemma 3.5.
@@
Let $T=2^{t}$ and $\beta=\frac{1}{\sqrt{2}}$. By Lemma 3.4, there exist states $|\psi_{0}\rangle,|\psi_{1}\rangle,\ldots,|\psi_{T-1}\rangle$ and Boolean phase oracles $B_{1}^{(k)},B_{2}^{(k)}$ for $10\leq k\leq T-$ 1, such that $|\psi_{k}\rangle=B_{2}^{(k)}{H}^{\otimes n}B_{1}^{(k)}{H}^{\otimes n}|0^{n}\rangle$ and
@@ 4 Lower bounds
In this section we prove ${T}$-count lower bounds matching the upper bounds in Theorem 1.1 and Theorem 1.2. Our lower bounds apply in a very general model of adaptive Clifford+${T}$ circuits that we now describe.
Adaptive Clifford+${T}$ Circuits.
Consider a quantum computation $\mathcal{A}$ with an $n$-qubit input register as well as an $a$-qubit ancilla register that is initialized in the all-zero state. The computation proceeds via a sequence of Clifford gates, ${T}$ gates, and single-qubit measurements in the computational basis. These operations may act on any of the $n+a$ qubits. Moreover, each Clifford or ${T}$ gate may be classically controlled on the outcomes of measurements that have occurred previously.
Let $r\in\{0,1\}^{*}$ be a binary string that describes the measurement outcomes in order obtained during the course of the computation. [^15] For maximum generality, we do not assume that the number of single-qubit measurements is bounded. If we knew there were exactly $m$ measurements, then $r$ would be an $m$-bit string. For each $r$, there is a Kraus operator $K_{r}$ which describes the corresponding sequence of ${T}$ gates, Clifford gates, and single-qubit projectors $|r_{i}\rangle\langle r_{i}|$ for $i=1,2,\ldots$. The adaptive Clifford+${T}$ circuit implements a quantum channel which maps the $n$-qubit input state $\rho$ to an $(n+a)$-qubit output state:
@@
For example, in the simplest case where there are no measurements and the circuit simply applies an $(n+a)$-qubit Clifford+${T}$ unitary $U$, $\mathcal{A}(\rho)=U\left(\rho\otimes|0^{a}\rangle\langle 0^{a}|\right)U^{\dagger}$. As another example, if the circuit applies a unitary $U$ and then measures the first qubit, then we have $\mathcal{A}(\rho)=\sum_{r\in\{0,1\}}(|r\rangle\langle r|\otimes I)U\left(\rho\otimes|0^{a}\rangle\langle 0^{a}|\right)U^{\dagger}(|r\rangle\langle r|\otimes I)$.
@@
Let $\mathcal{T}_{r}$ be the number of ${T}$ gates used by the circuit, conditioned on measurement outcome $r$. The maximal number of ${T}$ gates used by $\mathcal{A}$ is
@@
We also consider the expected ${T}$-count of $\mathcal{A}$ starting with the all-zeros input state:
@@
Finally, we define the expected ${T}$-count of $\mathcal{A}$ for the worst-case input state as follows
@@
where $\rho$ is any $n$-qubit mixed state and $\sum_{r}p_{r}(\rho)\mathcal{T}_{r}$ is the expected number of ${T}$ gates used in $\mathcal{A}(\rho)$. For example, if the circuit is unitary and has no measurements, then all three of these measures equal the ${T}$-count of the unitary.
@@
Here we used the fact that, if $\mathcal{A}$ is an adaptive Clifford+${T}$ circuit with an $n$-qubit input register, then so is $\mathcal{A}^{\prime}=I\otimes\mathcal{A}$ where the first register can have any number of qubits $n^{\prime}\geq 0$ and the second register is $n$ qubits (in particular, $\mathcal{A}^{\prime}$ is just $\mathcal{A}$ applied to the second register).
In the following we will prove lower bounds on the expected ${T}$-count $\mathcal{T}^{0}(\mathcal{A})$ (for state preparation) and $\mathcal{T}(\mathcal{A})$ (for unitary synthesis). These in turn imply lower bounds on the maximal ${T}$-count due to Equation 46.
@@
where $\mathrm{Tr}_{\mathrm{anc}}$ denotes the partial trace over the $a$-qubit ancilla register. However it is not hard to see that we can work with the closeness condition Equation 47 without loss of generality. Indeed, if an adaptive Clifford+${T}$ circuit $\mathcal{A}$ satisfies Equation 48, then we can construct another adaptive Clifford+${T}$ circuit $\mathcal{B}$ that satisfies Equation 47 and with the same expected ${T}$-count $\mathcal{T}^{0}(\mathcal{A})=\mathcal{T}^{0}(\mathcal{B})$. In particular $\mathcal{B}$ is obtained by first applying $\mathcal{A}$, then measuring the ancilla register in the computational basis, then applying a Pauli gate on the ancilla register which is controlled on the measurement outcome $z\in\{0,1\}^{a}$, that maps $|z\rangle\rightarrow|0^{a}\rangle$. Then $\mathcal{B}(|0^{n}\rangle\langle 0^{n}|)=\mathrm{Tr}_{\mathrm{anc}}(\mathcal{A}(|0^{n}\rangle\langle 0^{n}|))\otimes|0^{a}\rangle\langle 0^{a}|$ and the closeness condition Equation 47 holds for $\mathcal{B}$.
@@
There is an $n$-qubit state $|\psi\rangle$ such that any adaptive Clifford+${T}$ circuit $\mathcal{A}$ that prepares $|\psi\rangle$ up to error $\varepsilon$ in trace distance satisfies $\mathcal{T}^{0}(\mathcal{A})=\Omega\left(\sqrt{2^{n}\log(1/\varepsilon)}+\log(1/\varepsilon)\right)$.
This theorem lower bounds the expected ${T}$-count $\mathcal{T}^{0}(\mathcal{A})$ of an adaptive Clifford+${T}$ circuit–this is the most general setting we are able to handle. Note that if $\mathcal{A}$ is in fact a unitary $U$ composed of Clifford gates and ${T}$ gates (with no measurement), then its output state is a pure state $\mathcal{A}(\rho)=U(\rho\otimes|0^{a}\rangle\langle 0^{a}|)U^{\dagger}$ where $a$ is the number of ancillas used. In particular, there is only one Kraus operator $K=U$ in Equation 41. In this case the number of ${T}$ gates in the circuit for $U$ is equal to $\mathcal{T}_{\mathrm{max}}(\mathcal{A})$ which is also equal to $\mathcal{T}(\mathcal{A})$.
@@
There is an $n$-qubit diagonal unitary $D$ such that any adaptive Clifford+${T}$ circuit $\mathcal{A}$ that implements $D$ up to error $\varepsilon$ in diamond distance satisfies $\mathcal{T}(\mathcal{A})=\Omega\left(\sqrt{2^{n}\log(1/\varepsilon)}+\log(1/\varepsilon)\right)$.
Note that in the case where $\mathcal{A}$ is a unitary Clifford+${T}$ circuit with no measurements, the diamond distance is equivalent to the usual distance with respect to the operator norm, up to a global phase.
Finally, we establish the following ${T}$-count lower bound for synthesis of general unitaries.
@@
There is an $n$-qubit unitary $U$ such that any adaptive Clifford+${T}$ circuit $\mathcal{A}$ that implements $U$ up to error $\varepsilon$ in diamond distance satisfies $\mathcal{T}(\mathcal{A})=\Omega\left(2^{n}\sqrt{\log(1/\varepsilon)}+\log(1/\varepsilon)\right)$.

### v2

@@ 3 Quantum state preparation with optimal T-count
Our overall state preparation approach is similar to the one by Rosenthal [40]. Motivated by the Aaronson-Kuperberg problem [2, 1], [40] focuses on the number of Boolean oracle calls needed for efficient state preparation. While the context is different, the T{T}-count of the algorithm presented in [40] is near optimal. To obtain our Theorem 1.1, we combine Rosenthal’s algorithm with the crucial missing pieces Theorem 1.2 and Theorem 1.3. Along the way, we also simplify and improve some of Rosenthal’s analysis: our Lemma 3.2 simplifies and improves [40, Lemma 3.2], which was implicit in [25]; our Lemma 3.5 improves the query complexity in [40, Theorem 4.2].
We already know from Theorem 1.2 how to implement an arbitary diagonal unitary using the claimed number of T{T} gates and ancillas. For the state preparation task, this lets us specialize without loss of generality to the case where the target state has real amplitudes: after preparing a target state with real entries, we can apply a diagonal unitary to apply any desired complex phases.
@@
Let $|\psi\rangle$ be an arbitrary $n$-qubit state with real amplitudes. There exists an $n$-qubit state $|\phi\rangle$ with real amplitudes such that $\left\langle\phi\middle|\psi\right\rangle\geq\frac{1}{\sqrt{2}}$ and $|\phi\rangle=B_{2}{H}^{\otimes n}B_{1}{H}^{\otimes n}|0^{n}\rangle$ for some $n$-qubit Boolean phase oracles $B_{1}$ and $B_{2}$ (i.e., diagonal unitaries in which all diagonal entries are $\pm 1$).
@@
Let $B$ be an $n$-qubit Boolean phase oracle. Then $B$ can be prepared exactly by a Clifford+T{T} circuit using $O\left(\sqrt{2^{n}}\right)$ T{T} gates and ancillary qubits. [^20] This construction uses $O(2^{n})$ Clifford gates.
As a result of Fact 3.3, we can prepare the approximation $|\phi\rangle$ from Lemma 3.2 using only $O\left(\sqrt{2^{n}}\right)$ T{T} gates and ancilla qubits. We see that Lemma 3.2 already establishes the special case of Theorem 1.1 where the error $\varepsilon$ is a sufficiently large constant.
@@
Using Theorem 1.2 to correct phases, we assume without loss of generality that the target state $|\psi\rangle$ has real amplitudes. Then it suffices to analyze the T{T}-count and ancilla count for the circuit in Lemma 3.5. To this end, we analyze the components separately, each of which appears constant number of times in the whole circuit in Lemma 3.5. Assume $\varepsilon\leq 1/2$.
  • The reflection $2\left|0^{n+t}\right\rangle\left\langle 0^{n+t}\right|-I$ used in $R_{1}$ is a Boolean phase oracle and can be constructed by Fact 3.3 with $O\left(\sqrt{2^{n+t}}\right)=O\left(\sqrt{2^{n}\log(1/\varepsilon)}\right)$ T{T} gates and ancillas. [^21] This reflection can be implemented with much lower cost. We use Lemma 2.1 as a loose upper bound since it does not affect the asymptotic bound in the end.
@@
  • The Hadamard layer used in $V$ is Clifford and requires no T{T} gates or ancilla.
  • The Boolean phase oracle on $\mathcal{A}\mathcal{B}$ used in $V$ can be again prepared using Fact 3.3 with $O\left(\sqrt{2^{n+t}}\right)=O\left(\sqrt{2^{n}\log(1/\varepsilon)}\right)$ T{T} gates and ancillas.
  • The layer of single-qubit gates on $\mathcal{A}$ used in $V$ can be synthesized by Theorem 1.3 since $\mathcal{A}$ only has $t=\log\log(1/\varepsilon)+O(1)$ qubits. This uses $O\left(\log(1/\varepsilon)\right)$ T{T} gates and ancillas by Theorem 1.3. The approximation error does not blow up since this is only used a constant number of times in total.
@@
To get the coarse approximation of Lemma 3.2, we take our target state and apply a random diagonal unitary with $\pm 1$ entries followed by a binary Fourier transform (i.e., the Hadamard matrix). This does a reasonable job of flattening the amplitudes of the state, by which we mean that most of the amplitudes are now of similar magnitude. We then apply a diagonal unitary to make all the phases +1+1, which gives a state that has constant overlap with the uniform superposition $|+^{n}\rangle$. To implement this strategy we use the standard Khintchine inequality.
@@
After flattening, the state becomes essentially a uniform superposition with different phases. Moreover, in the case of interest where $|\psi\rangle$ is assumed to have real amplitudes, these phases are simply $\pm 1$. Then we can correct the Boolean phases and reverse the flattening procedure to obtain an approximation of the original state $|\psi\rangle$.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
In light of Lemma 3.8, we set $\alpha=\frac{1}{\sqrt{2}}$ and design $U_{\phi}$ as the BHBHB{H}B{H} circuit from Lemma 3.2. Note that all states used in Lemma 3.8 have real amplitudes since we start with $|\psi\rangle$ that has real amplitudes. Therefore circuits $U_{\phi}$ are all well-defined. ∎
@@
In this Section we describe the circuit that prepares an $n$-qubit state with optimal T{T}-count, establishing Lemma 3.5.
@@
Let $T=2^{t}$ and $\beta=\frac{1}{\sqrt{2}}$. By Lemma 3.4, there exist states $|\psi_{0}\rangle,|\psi_{1}\rangle,\ldots,|\psi_{T-1}\rangle$ and Boolean phase oracles $B_{1}^{(k)},B_{2}^{(k)}$ for $0\leq k\leq T-1$, such that $|\psi_{k}\rangle=B_{2}^{(k)}{H}^{\otimes n}B_{1}^{(k)}{H}^{\otimes n}|0^{n}\rangle$ and
@@ 4 Lower bounds
In this section we prove T{T}-count lower bounds matching the upper bounds in Theorem 1.1 and Theorem 1.2. Our lower bounds apply in a very general model of adaptive Clifford+T{T} circuits that we now describe.
Adaptive Clifford+T{T} Circuits.
Consider a quantum computation $\mathcal{A}$ with an $n$-qubit input register as well as an $a$-qubit ancilla register that is initialized in the all-zero state. The computation proceeds via a sequence of Clifford gates, T{T} gates, and single-qubit measurements in the computational basis. These operations may act on any of the $n+a$ qubits. Moreover, each Clifford or T{T} gate may be classically controlled on the outcomes of measurements that have occurred previously.
Let $r\in\{0,1\}^{*}$ be a binary string that describes the measurement outcomes in order obtained during the course of the computation. [^23] For maximum generality, we do not assume that the number of single-qubit measurements is bounded. If we knew there were exactly $m$ measurements, then $r$ would be an $m$-bit string. For each $r$, there is a Kraus operator $K_{r}$ which describes the corresponding sequence of T{T} gates, Clifford gates, and single-qubit projectors $|r_{i}\rangle\langle r_{i}|$ for $i=1,2,\ldots$. The adaptive Clifford+T{T} circuit implements a quantum channel which maps the $n$-qubit input state $\rho$ to an $(n+a)$-qubit output state:
@@
For example, in the simplest case where there are no measurements and the circuit simply applies an $(n+a)$-qubit Clifford+T{T} unitary $U$, $\mathcal{A}(\rho)=U\left(\rho\otimes|0^{a}\rangle\langle 0^{a}|\right)U^{\dagger}$. As another example, if the circuit applies a unitary $U$ and then measures the first qubit, then we have $\mathcal{A}(\rho)=\sum_{r\in\{0,1\}}(|r\rangle\langle r|\otimes I)U\left(\rho\otimes|0^{a}\rangle\langle 0^{a}|\right)U^{\dagger}(|r\rangle\langle r|\otimes I)$.
@@
Let $\mathcal{T}_{r}$ be the number of T{T} gates used by the circuit, conditioned on measurement outcome $r$. The maximal number of T{T} gates used by $\mathcal{A}$ is
@@
We also consider the expected T{T}-count of $\mathcal{A}$ starting with the all-zeros input state:
@@
Finally, we define the expected T{T}-count of $\mathcal{A}$ for the worst-case input state as follows
@@
where $\rho$ is any $n$-qubit mixed state and $\sum_{r}p_{r}(\rho)\mathcal{T}_{r}$ is the expected number of T{T} gates used in $\mathcal{A}(\rho)$. For example, if the circuit is unitary and has no measurements, then all three of these measures equal the T{T}-count of the unitary.
@@
Here we used the fact that, if $\mathcal{A}$ is an adaptive Clifford+T{T} circuit with an $n$-qubit input register, then so is $\mathcal{A}^{\prime}=I\otimes\mathcal{A}$ where the first register can have any number of qubits $n^{\prime}\geq 0$ and the second register is $n$ qubits (in particular, $\mathcal{A}^{\prime}$ is just $\mathcal{A}$ applied to the second register).
In the following we will prove lower bounds on the expected T{T}-count $\mathcal{T}^{0}(\mathcal{A})$ (for state preparation) and $\mathcal{T}(\mathcal{A})$ (for unitary synthesis). These in turn imply lower bounds on the maximal T{T}-count due to Equation 46.
@@
where $\mathrm{Tr}_{\mathrm{anc}}$ denotes the partial trace over the $a$-qubit ancilla register. However it is not hard to see that we can work with the closeness condition Equation 47 without loss of generality. Indeed, if an adaptive Clifford+T{T} circuit $\mathcal{A}$ satisfies Equation 48, then we can construct another adaptive Clifford+T{T} circuit $\mathcal{B}$ that satisfies Equation 47 and with the same expected T{T}-count $\mathcal{T}^{0}(\mathcal{A})=\mathcal{T}^{0}(\mathcal{B})$. In particular $\mathcal{B}$ is obtained by first applying $\mathcal{A}$, then measuring the ancilla register in the computational basis, then applying a Pauli gate on the ancilla register which is controlled on the measurement outcome $z\in\{0,1\}^{a}$, that maps $|z\rangle\rightarrow|0^{a}\rangle$. Then $\mathcal{B}(|0^{n}\rangle\langle 0^{n}|)=\mathrm{Tr}_{\mathrm{anc}}(\mathcal{A}(|0^{n}\rangle\langle 0^{n}|))\otimes|0^{a}\rangle\langle 0^{a}|$ and the closeness condition Equation 47 holds for $\mathcal{B}$.
@@
There is an $n$-qubit state $|\psi\rangle$ such that any adaptive Clifford+T{T} circuit $\mathcal{A}$ that prepares $|\psi\rangle$ up to error $\varepsilon$ in trace distance satisfies $\mathcal{T}^{0}(\mathcal{A})=\Omega\left(\sqrt{2^{n}\log(1/\varepsilon)}+\log(1/\varepsilon)\right)$.
This theorem lower bounds the expected T{T}-count $\mathcal{T}^{0}(\mathcal{A})$ of an adaptive Clifford+T{T} circuit–this is the most general setting we are able to handle. Note that if $\mathcal{A}$ is in fact a unitary $U$ composed of Clifford gates and T{T} gates (with no measurement), then its output state is a pure state $\mathcal{A}(\rho)=U(\rho\otimes|0^{a}\rangle\langle 0^{a}|)U^{\dagger}$ where $a$ is the number of ancillas used. In particular, there is only one Kraus operator $K=U$ in Equation 41. In this case the number of T{T} gates in the circuit for $U$ is equal to $\mathcal{T}_{\mathrm{max}}(\mathcal{A})$ which is also equal to $\mathcal{T}(\mathcal{A})$.
@@
There is an $n$-qubit diagonal unitary $D$ such that any adaptive Clifford+T{T} circuit $\mathcal{A}$ that implements $D$ up to error $\varepsilon$ in diamond distance satisfies $\mathcal{T}(\mathcal{A})=\Omega\left(\sqrt{2^{n}\log(1/\varepsilon)}+\log(1/\varepsilon)\right)$.
Note that in the case where $\mathcal{A}$ is a unitary Clifford+T{T} circuit with no measurements, the diamond distance is equivalent to the usual distance with respect to the operator norm, up to a global phase.
Finally, we establish the following T{T}-count lower bound for synthesis of general unitaries.
@@
There is an $n$-qubit unitary $U$ such that any adaptive Clifford+T{T} circuit $\mathcal{A}$ that implements $U$ up to error $\varepsilon$ in diamond distance satisfies $\mathcal{T}(\mathcal{A})=\Omega\left(2^{n}\sqrt{\log(1/\varepsilon)}+\log(1/\varepsilon)\right)$.